OpenStudy (anonymous):

Fan and Medal!!! What is the third term of a sequence that has a first term of 3 and a common ratio of 3? 1/3 9 27

OpenStudy (welshfella):

third term = a1 r^2 where a1 = first term , r = common ratio

OpenStudy (anonymous):

3*3^2 =9*3=27

OpenStudy (anonymous):

@welshfella is that right?

OpenStudy (michele_laino):

hint by definition of geometric sequence, we have: \[\Large \begin{gathered} {a_1} = 3 \hfill \\ {a_2} = 3{a_1} = ... \hfill \\ {a_3} = 3{a_2} = ... \hfill \\ \end{gathered} \]

OpenStudy (welshfella):

yes the nth term = a1 r^(n-1) here a1 = 3 and r = 3